# Regularized Conditional Diffusion Model for Multi-Task Preference Alignment

### v1

@@ Abstract
Sequential decision-making is desired to align with human intents and exhibit versatility across various tasks. Previous methods formulate it as a conditional generation process, utilizing return-conditioned diffusion models to directly model trajectory distributions. Nevertheless, the return-conditioned paradigm relies on pre-defined reward functions, facing challenges when applied in multi-task settings characterized by varying reward functions (versatility) and showing limited controllability concerning human preferences (alignment). In this work, we adopt multi-task preferences as a unified condition for both single- and multi-task decision-making, and propose preference representations aligned with preference labels. The learned representations are used to guide the conditional generation process of diffusion models, and we introduce an auxiliary objective to maximize the mutual information between representations and corresponding generated trajectories, improving alignment between trajectories and preferences. Extensive experiments in D4RL and Meta-World demonstrate that our method presents favorable performance in single- and multi-task scenarios, and exhibits superior alignment with preferences.
@@ 1 Introduction
In sequential decision-making, agents are trained to accomplish fixed or varied human-designed goals by interacting with the environment or learning from offline data. Two key objectives emerge during training: alignment, wherein agents are expected to take actions conforming to human intents expressed as crafted rewards or preferences, and versatility, denoting the capacity to tackle multiple tasks and generalize to unseen tasks. A promising avenue involves framing sequential decision-making as a sequence modeling problem via transformer Janner et al. (2021) or diffusion models Ho et al. (2020); Janner et al. (2022). This paradigm uses expressive models to capture the trajectory distributions and prevents unstable value estimation in conventional offline Reinforcement Learning (RL) methods Kostrikov et al. (2022). Particularly, utilizing return-conditioned or value-guided diffusion models to perform planning or trajectory generation achieves favorable performance in offline RL Ajay et al. (2023); Dong et al. (2023) and the multi-task variant Ni et al. (2023); He et al. (2023).
@@
Despite the great progress, applying diffusion models to sequential decision-making still faces challenges. (i) The condition generation process relies on a pre-defined reward function to provide return conditions. However, developing multiple task-specific reward functions in multi-task settings requires significant efforts Christiano et al. (2017) and may cause unintended behaviors Hejna III and Sadigh (2023). (ii) The condition generation process with classifier-free guidance Ho and Salimans (2021) often fails to ensure consistency between conditions and generations. As an example in Figure 1, the return-conditioned Decision Diffuser Ajay et al. (2023) struggles to achieve effective alignment for generated trajectories. Inspired by recent works Christiano et al. (2017); Ouyang et al. (2022); Go et al. (2023), we find that preferences offer more versatile supervision across multi-tasks than scalar rewards. Specifically, the trajectories from the $i$-th task are preferred over the $j$-th task if we set the $i$-th task as the target
@@
Figure 1: Illustration of the return-conditioned generation of Decision Diffuser Ajay et al. (2023) in hopper-medium-expert task. The return-conditional generation fails to align the returns and trajectories. We also show the desired alignment between the return conditions and the true returns of generated trajectories.
@@
task. Conversely, we can reverse the preference label when setting the target task to the $j$-th task. Therefore, we adopt preferences rather than rewards to guide the conditional diffusion models in offline multi-task RL. The generated trajectories are intended to align with preferences, prioritizing higher returns or specific tasks.
@@
To establish aligned and versatile conditional generation, our proposition involves adopting multi-task preferences and constructing a unified preference representation for both single- and multi-task scenarios, instead of learning scalar rewards from preference data Kim et al. (2023); Shin et al. (2023). The acquired representations are aligned with the preference labels. The representations subsequently serve as conditions to guide the conditional generation of diffusion models. In this case, two key challenges arise: (i) aligning the representations with preferences, and (ii) aligning the generated trajectories with representation conditions.
@@
To address the above challenges, we introduce Conditional Alignment via Multi-task Preference representations (CAMP) for multi-task preference learning. Specifically, we define multi-task preferences and extract preference representations from trajectory segments. (i) To align the representation with preferences, we propose a triplet loss and a Kullback-Leibler (KL) divergence loss to enable preferred and undesirable trajectories mapped into distinct zones in the representation space. The learned representations not only differentiate trajectories yielding higher returns but also discern trajectories across various tasks. Additionally, we learn an ‘optimal’ representation for each task to represent the trajectory with the highest preference. (ii) To align the generated trajectories with representation conditions in diffusion models, we introduce a Mutual Information (MI) regularization method. It augments representation-conditioned diffusion models with an auxiliary MI optimization objective, maximizing the correlation between the conditions and the generated outputs. During the inference for a specific task, we provide the ‘optimal’ representation for that task as the condition of the diffusion model, allowing the generated trajectories to adhere to desired preferences. Experiments on D4RL Fu et al. (2020) and Meta-World Yu et al. (2020) demonstrate the superiority of our method and the aligned performance in both multi-task and single-task scenarios.
@@ 2.1 Diffusion Models for MDP
Diffusion models have emerged as effective generative models adept at learning high-dimensional data distribution $p(x)$ Ho et al. (2020); Luo (2022); Oko et al. (2023). A diffusion model comprises a pre-defined forward process $q(x_{k}|x_{k-1})=\mathcal{N}(x_{k};\sqrt{\alpha_{k}}x_{k-1},(1-\alpha_{k})I)$ and a trainable reverse process $p_{\theta}(x_{k-1}|x_{k})$, where $k\in[1,K]$ denotes the timestep index, and $\alpha_{k}\in\mathbb{R}$ decides the variance schedule. By sampling Gaussian noise from $p(x_{K})$ and iteratively employing the denoising step $p_{\theta}(x_{k-1}|x_{k})$ for $K$ steps, diffusion models can generate $x_{0}\sim p(x)$. Moreover, if additional conditions $c$ are introduced into the denoising process such that $x_{k-1}\sim p_{\theta}(x_{k-1}|x_{k},c)$, diffusion models can also estimate the conditional distribution $p(x|c)$.
The Markov Decision Process (MDP) is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $P$ is the transition function, $r$ is the reward function, and $\gamma$ is a discount factor. We consider an offline setting where the policy is learned from a given dataset $\mathcal{D}$. For multi-task learning, we introduce an additional task space $\mathcal{T}$ that contains $m$ discrete tasks. Diffusion models have been used for offline RL to overcome the distribution shift caused by Temporal-Difference (TD) learning Kostrikov et al. (2022). Specifically, a diffusion model can formulate sequential decision-making as a conditional generative modeling problem by considering $x_{0}$ as a state sequence $(s_{t},\ldots,s_{t+h})$ for planning Ajay et al. (2023). The condition $c$ typically encompasses the return along the trajectory and is designed to generate trajectories with higher returns.
@@
where $q_{\psi}\coloneqq q_{\psi}(c|x_{0})$ represents the approximate variational posterior mapping $x_{0}$ to the condition $c$, and $q_{x_{k-1}}=q(x_{k-1}|x_{k},x_{0})$. The complete derivation is provided in §A.1. In practice, this optimization problem can be addressed via a score-matching objective Ho et al. (2020); Ho and Salimans (2021) as,
@@ 2.2 Preference Learning for Conditional Generation
In decision-making, human preferences are often applied on trajectory segments $\tau=[s_{i},a_{i}]_{i\in[1,h]}$ Wirth et al. (2017). For a trajectory pair $(\tau_{1},\tau_{2})$, human feedback yields a preference label $y\in\{0,1,0.5\}$ that indicates which segment a human prefers. Here, $y=1$ signifies $\tau_{1}\succ\tau_{2}$, $y=0$ signifies $\tau_{1}\prec\tau_{2}$, and $y=0.5$ means that two trajectories have the same preference. Previous studies commonly employ the Bradley-Terry (BT) model Bradley and Terry (1952) to derive a reward function $\hat{r}$ from such preferences. Considering learning with offline preference data Shin et al. (2023), and given a dataset $\mathcal{D}_{\tau}=\{(\tau_{1},\tau_{2},y)\}$, $\hat{r}$ is learned by maximizing the following objective:
@@
In what follows, we simplify the notation by omitting the label $y$ and denote the preference data as $\{(\tau^{+},\tau^{-})\}$, where we have $\tau^{+}\succ\tau^{-}$. Previous methods Lee et al. (2021); Kim et al. (2023) based on the BT model follow a two-phase learning process, first deriving a reward function from preference data and then training a policy with RL. Nevertheless, this process hinges on the assumption that pairwise preferences can be substituted with a reward model that can generalize to out-of-distribution data Azar et al. (2023). Moreover, these methods often require complex reward models Kim et al. (2023) when preferences are intricate Coste et al. (2023); Hejna et al. (2023).
@@ 3.2 Representation Aligned Trajectory Generation
Given preference-based versatile representations, we train a diffusion model to generate trajectories that align with the representations. Prior work utilizes classifier-free guidance and aligns the generated samples with low-dimensional conditions, such as returns. However, we find this is insufficient to capture the complex relationship between conditions and trajectories. Using the representative method Decision Diffuser Ajay et al. (2023) as an example, Figure 1 reveals the relationship between the return conditions and true returns of generated trajectories in the Hopper-medium-expert task. While it is desirable to generate trajectories with higher returns as we increase the return condition, in practice, the attributes of generated samples do not exhibit a smooth transition with changes in the condition. Empirically, Decision Diffuser uses a target return of 0.8-0.9 for the generation process. This difficulty becomes more severe in our method because we cannot exclusively search the high-dimensional representation space and find an empirically effective representation. Therefore, it is critical to enhance the alignment between generated trajectories and previously learned preference representations.
@@
Inspired by recent works on generative models Zhao et al. (2017); Chen et al. (2016); Wang et al. (2023a), we introduce an auxiliary optimization objective aimed at strengthening the association between representations $w\sim q_{\psi}(w|\tau)$ and the generated trajectories $\tau_{0}$. Specifically, we augment the learning objective from Eq. (2) with a regularization term based on mutual information between $\tau_{0}$ and $w$. Formally, we train the conditional diffusion model using the following combined objectives:
@@
In practice, sampling $\tau_{0}$ from $p(\tau_{0})$ to maximize the MI objective requires the diffusion model to denoise $K$ steps from $\tau_{K}$. This process imposes huge computational burden and may lead to potential memory overflow due to the gradient propagation across multi-steps. Consequently, there arises a need for an approximate objective, and an alternative approach is to replace the optimization on $I(\tau_{0};w)$ with $I(\tau_{k};w)$. This substitution is motivated by considering the relationship between $\tau_{0}$ and $\tau_{k}$, as described by the diffusion process $\tau_{k}=\sqrt{\bar{\alpha}}_{t}\tau_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{0}\coloneqq f_{\rm diffuse}(\tau_{0})$. According to the data processing inequality Beaudry and Renner (2012), we have
@@
[…]
The first term resembles the ELBO term in Eq. (1) denoted as $\mathcal{L}_{\rm elbo}(x_{0}=\tau_{0},c=w)$, which is the same as standard conditional diffusion models in Decision Diffuser Ajay et al. (2023). The ELBO term aims to estimate the trajectory distribution via a conditional diffusion process, thus we adopt a similar conditional score-matching objective to optimize it. The second term contributes to enhancing the alignment between $\tau_{0}$ and $w$, where $p_{\psi}(w)$ is empirically averaged on samples $\tau\in\mathcal{D}$ via $f_{\psi}$, and $p(\tau_{k})\propto f_{\text{diffuse}}(\tau_{0})$ with $\tau_{0}$ sampling from $q(\tau_{0})$. We can minimize the KL divergence to optimize $q_{\phi}(w|\tau_{k})$, a variational predictor to predict the condition $w$ from the denoised sample $\tau_{k}$. In practice, we instantiate it with a neural network represented by $\phi$, taking predicted noises $\epsilon_{\theta}(\tau_{k})$ as inputs.
@@ 4 Related Work
Diffusion models exhibit significant advantages in generative modeling and characterizing complex distributions. On the one hand, they can be applied for modeling policies, with many research works Wang et al. (2023b); Chi et al. (2023); Hansen-Estruch et al. (2023); Chen et al. (2023); Pearce et al. (2023) suggesting that diffusion models are more effective at capturing multi-modal action distributions than other policy types like energy-based policies. On the other hand, diffusion models are adept at directly predicting trajectories via conditional generation, adopting conditions such as value gradients Janner et al. (2022); Liang et al. (2023); Ni et al. (2023), returns Ajay et al. (2023), or prompts He et al. (2023). Our work also trains diffusion models to generate trajectories, but focuses on guidance from preference representations. Regarding alignment with human preferences, recent works include the integration with designed attributes for various behaviors Dong et al. (2023), and preference augmentation techniques to improve trajectory generation Anonymous (2024). While these works focus on single-task settings, our work seeks a versatile solution by considering multi-task preferences.
@@
Current methods of learning from preferences can be categorized into two groups: direct learning and indirect learning. Indirect learning methods involve learning a reward model and incorporating existing RL algorithms. They employ techniques like off-policy learning Lee et al. (2021), pseudo-labeling and data augmentation Park et al. (2022), iterative updates of reward models and policies Liu et al. (2022), or leveraging Transformer architectures to learn reward models Kim et al. (2023). On the other hand, direct learning methods bypass reward model learning and directly incorporate human preferences into the learning process. This approach involves various strategies, such as combining decision transformer styles with preference embeddings Kang et al. (2023), mapping reward functions to optimal policies Rafailov et al. (2023); Zhou et al. (2023), or aligning models using extended Bradley-Terry comparisons Song et al. (2023). This work falls under the category of offline direct learning approaches, framing it as a sequence modeling problem.
Our work considers challenging multi-task settings. While much prior research attempts to find Pareto optimal solutions considering the trade-offs between different preferences, these methods often require heuristic vector selection for unknown Pareto fronts Mahapatra and Rajan (2020); Lin et al. (2019), millions of pre-collected preference labels and further online queries Hejna III and Sadigh (2023), or combination with Gaussian processes to learn preference relations Birlutiu et al. (2009). In contrast, our approach does not require any heuristic methods and learns trajectory representations aligned with preferences from offline data.

### v2

@@ Abstract
Sequential decision-making can be formulated as a conditional generation process, with targets for alignment with human intents and versatility across various tasks. Previous return-conditioned diffusion models manifest comparable performance but rely on well-defined reward functions, which requires amounts of human efforts and faces challenges in multi-task settings. Preferences serve as an alternative but recent work rarely considers preference learning given multiple tasks. To facilitate the alignment and versatility in multi-task preference learning, we adopt multi-task preferences as a unified framework. In this work, we propose to learn preference representations aligned with preference labels, which are then used as conditions to guide the conditional generation process of diffusion models. The traditional classifier-free guidance paradigm suffers from the inconsistency between the conditions and generated trajectories. We thus introduce an auxiliary regularization objective to maximize the mutual information between conditions and corresponding generated trajectories, improving their alignment with preferences. Experiments in D4RL and Meta-World demonstrate the effectiveness and favorable performance of our method in single- and multi-task scenarios.
@@ 1 Introduction
In sequential decision-making, agents are trained to accomplish fixed or varied human-designed goals by interacting with the environment or learning from offline data. Two key objectives emerge during training: alignment, wherein agents are expected to take actions conforming to human intents expressed as crafted rewards or preferences, and versatility, denoting the capacity to tackle multiple tasks and generalize to unseen tasks. A promising avenue involves framing sequential decision-making as a sequence modeling problem via transformer [1] or diffusion models [2, 3]. This paradigm uses expressive models to capture the trajectory distributions and prevents unstable value estimation in conventional offline Reinforcement Learning (RL) methods [4]. Particularly, utilizing return-conditioned or value-guided diffusion models to perform planning or trajectory generation achieves favorable performance in offline RL [5, 6] and the multi-task variant [7, 8].
@@
Despite the great progress, applying diffusion models to sequential decision-making still faces challenges. (i) The condition generation process relies on a pre-defined reward function to provide return conditions. However, developing multiple task-specific reward functions in multi-task settings requires significant efforts [9] and may cause unintended behaviors [10]. (ii) The condition generation process with classifier-free guidance [11] often fails to ensure consistency between conditions and generations. As an example in Figure 1, the return-conditioned Decision Diffuser [5] struggles to achieve effective alignment for generated trajectories. Inspired by recent works [9, 12, 13], we find that preferences offer more versatile supervision across multi-tasks than scalar rewards. Specifically, the trajectories from the $i$-th task are preferred
@@
Figure 1: Illustration of the return-conditioned generation of Decision Diffuser [5] in hopper-medium-expert task. Existing return-conditional diffusion models fail to align generated trajectories with the return condition, while the red line indicates the desired relationship between the return conditions and true returns of generated trajectories [14].
@@
over the $j$-th task if we set the $i$-th task as the target task. Conversely, we can reverse the preference label when setting the target task to the $j$-th task. Therefore, we adopt preferences rather than rewards to guide the conditional diffusion models in offline multi-task RL. The generated trajectories are intended to align with preferences, prioritizing higher returns or specific tasks.
@@
To establish aligned and versatile conditional generation, our proposition involves adopting multi-task preferences and constructing a unified preference representation for both single- and multi-task scenarios, instead of learning scalar rewards from preference data [15, 16]. The acquired representations are aligned with the preference labels. The representations subsequently serve as conditions to guide the conditional generation of diffusion models. In this case, two key challenges arise: (i) aligning the representations with preferences, and (ii) aligning the generated trajectories with representation conditions.
@@
To address the above challenges, we introduce Conditional Alignment via Multi-task Preference representations (CAMP) for multi-task preference learning. Specifically, we define multi-task preferences and extract preference representations from trajectory segments. (i) To align the representation with preferences, we propose a triplet loss and a Kullback-Leibler (KL) divergence loss to enable preferred and undesirable trajectories mapped into distinct zones in the representation space. The learned representations not only differentiate trajectories yielding higher returns but also discern trajectories across various tasks. Additionally, we learn an ‘optimal’ representation for each task to represent the trajectory with the highest preference. (ii) To align the generated trajectories with representation conditions in diffusion models, we introduce a Mutual Information (MI) regularization method. It augments representation-conditioned diffusion models with an auxiliary MI optimization objective, maximizing the correlation between the conditions and the generated outputs. During the inference for a specific task, we provide the ‘optimal’ representation for that task as the condition of the diffusion model, allowing the generated trajectories to adhere to desired preferences. Experiments on D4RL [17] and Meta-World [18] demonstrate the superiority of our method and the aligned performance in both multi-task and single-task scenarios.
@@ 2.1 Diffusion Models for MDP
Diffusion models have emerged as effective generative models adept at learning high-dimensional data distribution $p(x)$ [2, 19, 20]. A diffusion model comprises a pre-defined forward process $q(x_{k}|x_{k-1})=\mathcal{N}(x_{k};\sqrt{\alpha_{k}}x_{k-1},(1-\alpha_{k})I)$ and a trainable reverse process $p_{\theta}(x_{k-1}|x_{k})$, where $k\in[1,K]$ denotes the timestep index, and $\alpha_{k}\in\mathbb{R}$ decides the variance schedule. By sampling Gaussian noise from $p(x_{K})$ and iteratively employing the denoising step $p_{\theta}(x_{k-1}|x_{k})$ for $K$ steps, diffusion models can generate $x_{0}\sim p(x)$. Moreover, if additional conditions $c$ are introduced into the denoising process such that $x_{k-1}\sim p_{\theta}(x_{k-1}|x_{k},c)$, diffusion models can also estimate the conditional distribution $p(x|c)$.
The Markov Decision Process (MDP) is defined by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $P$ is the transition function, $r$ is the reward function, and $\gamma$ is a discount factor. We consider an offline setting where the policy is learned from a given dataset $\mathcal{D}$. For multi-task learning, we introduce an additional task space $\mathcal{T}$ that contains $m$ discrete tasks. Diffusion models have been used for offline RL to overcome the distribution shift caused by Temporal-Difference (TD) learning [4]. Specifically, a diffusion model can formulate sequential decision-making as a conditional generative modeling problem by considering $x_{0}$ as a state sequence $(s_{t},\ldots,s_{t+h})$ for planning [5]. The condition $c$ typically encompasses the return along the trajectory and is designed to generate trajectories with higher returns.
@@
where $q_{\psi}\coloneqq q_{\psi}(c|x_{0})$ represents the approximate variational posterior mapping $x_{0}$ to the condition $c$, and $q_{x_{k-1}}=q(x_{k-1}|x_{k},x_{0})$. The complete derivation is provided in §A.1. In practice, this optimization problem can be addressed via a score-matching objective [2, 11] as,
@@ 2.2 Preference Learning for Conditional Generation
In decision-making, human preferences are often applied on trajectory segments $\tau=[s_{i},a_{i}]_{i\in[1,h]}$ [22]. For a trajectory pair $(\tau_{1},\tau_{2})$, human feedback yields a preference label $y\in\{0,1,0.5\}$ that indicates which segment a human prefers. Here, $y=1$ signifies $\tau_{1}\succ\tau_{2}$, $y=0$ signifies $\tau_{1}\prec\tau_{2}$, and $y=0.5$ means that two trajectories have the same preference. Previous studies commonly employ the Bradley-Terry (BT) model [23] to derive a reward function $\hat{r}$ from such preferences. Considering learning with offline preference data [16], and given a dataset $\mathcal{D}_{\tau}=\{(\tau_{1},\tau_{2},y)\}$, $\hat{r}$ is learned by maximizing the following objective:
@@
In what follows, we simplify the notation by omitting the label $y$ and denote the preference data as $\{(\tau^{+},\tau^{-})\}$, where we have $\tau^{+}\succ\tau^{-}$. Previous methods [24, 15] based on the BT model follow a two-phase learning process, first deriving a reward function from preference data and then training a policy with RL. Nevertheless, this process hinges on the assumption that pairwise preferences can be substituted with a reward model that can generalize to out-of-distribution data [25]. Moreover, these methods often require complex reward models [15] when preferences are intricate [26, 27].
@@ 3.2 Representation Aligned Trajectory Generation
Given preference-based versatile representations, we train a diffusion model to generate trajectories that align with the representations. Prior work utilizes classifier-free guidance and aligns the generated samples with low-dimensional conditions, such as returns. However, we find this is insufficient to capture the complex relationship between conditions and trajectories. Using the representative method Decision Diffuser [5] as an example, Figure 1 reveals the relationship between the return conditions and true returns of generated trajectories in the Hopper-medium-expert task. While it is desirable to generate trajectories with higher returns as we increase the return condition, in practice, the attributes of generated samples do not exhibit a smooth transition with changes in the condition. Empirically, Decision Diffuser uses a target return of 0.8-0.9 for the generation process. This difficulty becomes more severe in our method because we cannot exclusively search the high-dimensional representation space and find an empirically effective representation. Therefore, it is critical to enhance the alignment between generated trajectories and previously learned preference representations.
@@
Inspired by recent works on generative models [28, 29, 21], we introduce an auxiliary optimization objective aimed at strengthening the association between representations $w\sim q_{\psi}(w|\tau)$ and the generated trajectories $\tau_{0}$. Specifically, we augment the learning objective from Eq. (2) with a regularization term based on mutual information between $\tau_{0}$ and $w$. Formally, we train the conditional diffusion model using the following combined objectives:
@@
In practice, sampling $\tau_{0}$ from $p(\tau_{0})$ to maximize the MI objective requires the diffusion model to denoise $K$ steps from $\tau_{K}$. This process imposes huge computational burden and may lead to potential memory overflow due to the gradient propagation across multi-steps. Consequently, there arises a need for an approximate objective, and an alternative approach is to replace the optimization on $I(\tau_{0};w)$ with $I(\tau_{k};w)$. This substitution is motivated by considering the relationship between $\tau_{0}$ and $\tau_{k}$, as described by the diffusion process $\tau_{k}=\sqrt{\bar{\alpha}}_{t}\tau_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{0}\coloneqq f_{\rm diffuse}(\tau_{0})$. According to the data processing inequality [30], we have
@@
$\mathcal{L}_{\rm I}(\theta,\phi)=\mathbb{E}_{q(\tau_{0})}[\mathcal{L}_{\rm elbo}(\tau_{0},w)]-\zeta\cdot\mathbb{E}_{p(\tau_{k})}\left[D_{\text{KL}}(p_{\psi}(w))\|q_{\phi}(w|\tau_{k})\right].$ (8)
The first term resembles the ELBO term in Eq. (1) denoted as $\mathcal{L}_{\rm elbo}(x_{0}=\tau_{0},c=w)$, which is the same as standard conditional diffusion models in Decision Diffuser [5]. The ELBO term aims to estimate the trajectory distribution via a conditional diffusion process, thus we adopt a similar conditional score-matching objective to optimize it. The second term contributes to enhancing the alignment between $\tau_{0}$ and $w$, where $p_{\psi}(w)$ is empirically averaged on samples $\tau\in\mathcal{D}$ via $f_{\psi}$, and $p(\tau_{k})\propto f_{\text{diffuse}}(\tau_{0})$ with $\tau_{0}$ sampling from $q(\tau_{0})$. We can minimize the KL divergence to optimize $q_{\phi}(w|\tau_{k})$, a variational predictor to predict the condition $w$ from the denoised sample $\tau_{k}$. In practice, we instantiate it with a neural network represented by $\phi$, taking predicted noises $\epsilon_{\theta}(\tau_{k})$ as inputs.
@@ 4 Related Work
Diffusion models exhibit significant advantages in generative modeling and characterizing complex distributions. On the one hand, they can be applied for modeling policies, with many research works [31, 32, 33, 34, 35, 36] suggesting that diffusion models are more effective at capturing multi-modal action distributions than other policy types like energy-based policies. On the other hand, diffusion models are adept at directly predicting trajectories via conditional generation, adopting conditions such as value gradients [3, 37, 7], returns [5], or prompts [8]. Our work also trains diffusion models to generate trajectories, but focuses on guidance from preference representations. Regarding alignment with human preferences, recent works include the integration with designed attributes for various behaviors [6], fine-tuned policy with preferences [38, 39], and preference augmentation techniques to improve trajectory generation [40]. While these works focus on single-task settings, our work seeks a versatile solution by considering multi-task preferences.
@@
Current methods of learning from preferences can be categorized into two groups: direct learning and indirect learning. Indirect learning methods involve learning a reward model and incorporating existing RL algorithms. They employ techniques like off-policy learning [24], pseudo-labeling and data augmentation [41], iterative updates of reward models and policies [42], or leveraging Transformer architectures to learn reward models [15]. On the other hand, direct learning methods bypass reward model learning and directly incorporate human preferences into the learning process. This approach involves various strategies, such as combining decision transformer styles with preference embeddings [43], mapping reward functions to optimal policies [44, 45], or aligning models using extended Bradley-Terry comparisons [46]. This work falls under the category of offline direct learning approaches, framing it as a sequence modeling problem.
Our work considers challenging multi-task settings. While much prior research attempts to find Pareto optimal solutions considering the trade-offs between different preferences, these methods often require heuristic vector selection for unknown Pareto fronts [47, 48], millions of pre-collected preference labels and further online queries [10], or combination with Gaussian processes to learn preference relations [49]. In contrast, our approach does not require any heuristic methods and learns trajectory representations aligned with preferences from offline data.